OpenStudy (lina777):

I will fan and medal!! What is tan(arcsin x/2)? Please explain this to me using a right triangle :)

OpenStudy (anthonyym):

Do you need to simplify it because I don't think you'll get a numerical answer

OpenStudy (lina777):

Yes, simplification :) @anthonyym

OpenStudy (sshayer):

\[\arcsin \frac{ x }{ 2 }=y,\frac{ x }{ 2 }=\sin y\] 

OpenStudy (sshayer):

\[\tan \left( arc \sin^{-1} \frac{ x }{ 2 } \right)=\tan y=\frac{ x }{ \sqrt{4-x^2} }\]